OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 5 2x - 4y = 10 Infinitely Many Solutions No Solutions (0, 0) (5, 10)

OpenStudy (anonymous):

Multiply the first equation by 2. You'll see that these two equations are actually the same, so there is no way to isolate each variable. You can pick arbitrary values for one of the variables and solve for the other variable. Infinitely many.

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3) so for this one wouldn't it be B or D?

OpenStudy (anonymous):

2,3

OpenStudy (anonymous):

It's not really a guessing game :) You can figure this out! y=11-4x Substitute this into the second eq: x+2(11-4x)=8 x+22-8x=8 -7x=-14 7x=14 take it from here?

OpenStudy (anonymous):

alright its 2,3